# Adaptive Super-Resolution Imaging Without Prior Knowledge Using a Programmable Spatial-Mode Sorter

### v1

@@ 4.2 First Stage
In the first stage of this experiment, we acquired direct imaging data by sweeping the point source (20x20 pixels on the SLM) across the full field of view (200 pixels vertically) 5 times. The reason that we did it 5 times is to be able to combine images from different sweeps which will supposedly make it more difficult to estimate both the centroid and the separation and will make the scenario more realistic. Once all the data was acquired, we randomly chose 200 different centroids and for each random centroid we chose the appropriate frame from a random data set (each one of the 5 point-source sweeps is considered one data set). After the chosen frames were combined into one image (an example of that with all tested separations is shown in Fig.4), we estimated the centroid using the center of mass method.
After the centroid was estimated, it was fed into to the MPLC based, phase mask generation algorithm which created the appropriate phase masks to sort the wanted modes where the estimated centroid is set to be the center of the scene. A total of 2400 (12 separations and 200 data points for each separation) centroids and phase mask sets were generated.
@@ 4.3 Second Stage
The second stage of the experiment focused on acquiring the mode sorting data (shown in Fig.5). In this stage, we imaged the point sources (with the mode sorter) at a specified location, which is the same randomly chosen location discussed as part of the first stage of the experiment using the appropriate phase masks set that was generated using the centroid estimation from the matching direct imaging data. This was done 4800 times (each data point required 2 point sources, and we did it 200 times for the 12 separations that were tested).
@@ 5.1 Centroid Estimation
$\phi_{x}=\frac{\sum_{i=-256}^{i=255}\sum_{j=-256}^{j=255}M_{x_{i,j}}\times Im_{i,j}}{\sum_{i=-256}^{i=255}\sum_{j=-256}^{j=255}Im_{i,j}},\rm{and}$ (2a)
$\phi_{y}=\frac{\sum_{i=-256}^{i=255}\sum_{j=-256}^{j=255}M_{y_{i,j}}\times Im_{i,j}}{\sum_{i=-256}^{i=255}\sum_{j=-256}^{j=255}Im_{i,j}}.$ (2b)
This method is valid in our case since we used equally intense point sources, however, this approach will need to be modified if point sources of different intensities are under test.
@@ 5.2 Separation Estimation- Direct Imaging
Since in our experiment we used a finite circular hard aperture we can describe the intensity pattern resulting from a point source as an airy disk:
$S(x,y)=\left(\frac{2\times J\left(\frac{\pi\times\sqrt{x^{2}+y^{2}}\times D_{\rm{ph}}}{\lambda\times f}\right)}{\frac{\pi\times\sqrt{x^{2}+y^{2}}\times D_{\rm{ph}}}{\lambda\times f}}\right)^{2}\rm{,}$ (3)
where $J(.)$ is the Bessel function, $D_{\rm{ph}}$ is the diameter of the circular aperture, $f$ is the focal length of the lens used in the system, and $\lambda$ is the wavelength.
Because in this experiment we have two point sources separated in 1D we can describe the intensity pattern on the detector as such:
$\Psi(x,y)=S(x+\phi_{x}+\frac{\theta}{2},y+\phi_{y})+S(x+\phi_{x}-\frac{\theta}{2},y+\phi_{y}).$ (4)
@@
$P_{D}(\{x_{i}\}_{i=1}^{n_{1}},\phi_{x}|\theta)=\prod_{i=1}^{n_{1}}\Psi(x_{i},y),$ (5)
where $\theta$ is the separation between the point source, $x_{i}$ stands for the position of the $i^{\text{th}}$ detected photon, and $n_{1}$ stands for the total number of photons detected in the direct imaging stage.
Now, by maximizing the likelihood function we get our estimation for the point-sources separation :
$\hat{\theta}_{\rm{direct}}=\mathrm{argmax}_{\theta}(P_{D}(\{x_{i}\}_{i=1}^{n_{1}},\phi_{x}|\theta)).$ (6)
@@ 5.3.1 Maximum Likelihood
In our experiment, we sorted Zernike or Airy disk modes. However, for simplicity, and due to the similarity of their plots in the super-resolution regime we matched our binary SPADE photon counting results to the probability of a photon landing in the 0th and 1st Hermite-Gauss modes as a function of the separation, $\theta$, which can be described as such:
$\psi_{0}=\frac{\exp{\frac{-\theta^{2}}{\sigma^{2}}}}{\max(\exp{\frac{-\theta^{2}}{\sigma^{2}}})},\rm{and}$ (7a)
$\psi_{1}\approx 1-\psi_{0},$ (7b)
where:
$\sigma=1.22\times\frac{\lambda\times f}{D_{\rm{ph}}}.$ (8)
$\psi_{1}$ can be described as it is in eq.7b because we are doing binary SPADE and therefore, all the photons that will not be counted in the $\psi_{0}$ spot, will be counted from the $\psi_{1}$ spot.
It is important to note that when applying this model to our experimental cross-talk ($C$) (see supplementary material section A for the cross-talk definition) was taken into account such that $\psi_{0}$ and $\psi_{1}$ become:
$\psi_{0}=C\times\frac{\exp{\frac{-\theta^{2}}{\sigma^{2}}}}{\max(\exp{\frac{-\theta^{2}}{\sigma^{2}}})},\rm{and}$ (9a)
$\psi_{1}=\exp{\frac{-\theta^{2}}{\sigma^{2}}}\times 2\times(\frac{\theta}{\sigma})^{2}.$ (9b)
@@
$P_{B}(I_{0},n_{2}|\theta)=\begin{pmatrix}n_{2}\\ I_{0}\end{pmatrix}(\psi_{0})^{I_{0}}(\psi_{1})^{n_{2}-I_{0}},$ (10)
@@
Now, to estimate the most likely $\theta$ we maximize the likelihood function $P_{B}$:
$\hat{\theta}_{\rm{Sorting}}=\mathrm{argmax}_{\theta}(P_{B}(I_{0},n_{2}|\theta)).$ (11)
@@ 5.3.2 Comparison to Characterization
The exact number of photons that gets detected at every data point will vary. Because of that, rather than comparing the number of photons in each mode to our characterization we found the ratio between $\psi_{0}$ and $\psi_{1}$ from our data and matched it to the same ratio from our interpolated system’s characterization (Fig. 3). By matching the ratios we could then find what is the separation that is the most fitting for the data.
@@ 5.4 Separation Estimation- 2-Stage
As mentioned in section 3A, when we try to estimate $\theta$ using our mode sorter, we still use direct imaging for about half of the photons for the centroid estimation. In this method instead of just using these photons for centroid estimation we also use them for the separation estimation. This leads to a new likelihood function that utilizes all the photons collected:
$P_{{\operatorname{2-stage}}}(\{x_{i}\}_{i=1}^{n_{1}},\phi_{x},I_{0},n_{2}|\theta)=P_{D}(\{x_{i}\}_{i=1}^{n_{1}},\phi_{x}|\theta)\times P_{B}(I_{0},n_{2}|\theta).$ (12)
Therefore, the estimated $\theta$ using this method can be found by maximizing the above likelihood function:
$\hat{\theta}_{{\operatorname{2-stage}}}=\mathrm{argmax}_{\theta}(P_{{\operatorname{2-stage}}}(\{x_{i}\}_{i=1}^{n_{1}},\phi_{x},I_{0},n_{2}|\theta)).$ (13)
@@ 6 Results
The goal of this experiment was to prove the superiority of our adaptive mode sorter in separation estimation below the Rayleigh limit when compared to traditional direct imaging.
@@
Figure 6: Monte-Carlo simulation that shows the mean squared error (MSE) of separation estimation using 10000 photons of direct imaging (blue circles), 5000 photons in direct imaging for centroid estimation, and 5000 photons for sorting (orange circles), and 5000 photons in direct imaging for centroid estimation and for separation estimation, and 5000 photons for sorting (yellow circles). Note that because in our simulation we assumed the absence of bias the plot for the MSE vs. the separation will look identical to the plot for variance vs. separation
@@
Figure 7: a) Mean squared error of the separation estimation , b) Variance of the separation estimation using 10000 photons of direct imaging (blue circles), 5000 photons in direct imaging for centroid estimation and for separation estimation,5000 photons for sorting with the maximum likelihood method (orange circles), and 5000 photons for sorting with the comparison to characterization method (purple circles)
@@
Before conducting the experiment, we ran a Monte Carlo simulation. The results of the simulation are promising as the separation estimation MSE in the deep sub-Rayleigh regime (below 0.5 Rayleigh units of separation) shows an improvement by a factor of up to 12.7. Note that in the Monte Carlo simulation, we assumed a perfect mode-sorter, with no crosstalk, (see supplementary information), a perfect detector (no noise). The simulation results can be seen in Fig. 6. In this figure, when looking at the simulation results when only using the sorting data for separation estimation (orange circles), one can notice a counter-intuitive phenomenon, the MSE increases with the separation. This cam be explained by the fact that at very small separations most of the Fisher information exists in the modes that we sort (Zernike00 and Zernike01) and as the separation increases there is more Fisher information that lives in higher order modes. Another way to think about it is by looking at fig.3. When the separation is very small, there are very few photons arriving at the Zernike01 spot and every change in that is easily detected. However, as the separation increases, small changes at the intensity of the Zernike01 spot become harder to detect and therefore making an error becomes more likely.
Fig.7a presents our experimental MSE results. One can see that in the deep sub-Rayleigh regime (up to 0.2 Rayleigh units of separation) our mode sorter is up to 2.56 times more accurate than experimental direct imaging and only 3.63 times less accurate than the Monte Carlo simulation which simulates a perfect detector.
In Fig.7b one can see that the variance of our mode sorter is lower than that of direct imaging up to 0.3 Rayleigh units of separation and is very close to it at 0.4 Rayleigh units of separation.
This tells us that there is an unexpected source of bias in our mode sorting data (this will be further discussed in the Supplementary Materials in section F). Some possible sources of error that may have degraded or biased our data are: crosstalk, although it was taken into account in the data processing, misalignment of the system, generally and with respect to the centroid, and shot noise in the detector. See supplementary information for further discussion of the sources of error.
7 Acknowledgments

### v2

@@ 4.2 First Stage
In the first stage of this experiment, we acquired direct imaging data by sweeping the point source (20x20 pixels on the SLM, limited by the amount of light needed in the system) across the full field of view (200 pixels vertically) 5 times. At this point it is important to recall that this was an emulated experiment, therefore, 5 sweeps allowed us to combine images from different sweeps to create our emulated, 2 point source scene. Combining images from different sweeps to construct our scene will make it more difficult to estimate both the centroid and the separation and will make this emulated scenario more realistic. Once all the data was acquired, we randomly chose 200 different centroids and for each random centroid we chose the appropriate frame from a random data set (each one of the 5 point-source sweeps is considered one data set). After the chosen frames were combined into one image (an example of that with all tested separations is shown in Fig.4), we estimated the centroid using the center of mass method.
After the centroid was estimated, it was fed into to the MPLC phase mask generation algorithm which created the appropriate phase masks to sort the wanted modes, where the estimated centroid is set to be the center of the scene for each computation. A total of 2400 (12 separations and 200 data points for each separation) centroids and phase mask sets were generated. While the computation time required to generate a set of phase masks for each trial necessitated that we operate our two-stage experiment asynchronously, in a future system a set of phase masks could be chosen from a pre-computed set based on the estimate from the preliminary measurement.
@@ 4.3 Second Stage
The second stage of the experiment focused on acquiring the mode sorting data (shown in Fig.5). In this stage, we computationally combined the images acquired from displaying the two point sources separately, where the data was the camera acquisitions with the MPLC active. This was done 4800 times (each data point required 2 point sources, and we did it 200 times for the 12 separations that were tested).
@@ 5.1 Centroid Estimation
$\hat{\phi}_{x}=\frac{\sum_{i=-256}^{i=255}\sum_{j=-256}^{j=255}M_{x_{i,j}}\times Im_{i,j}}{\sum_{i=-256}^{i=255}\sum_{j=-256}^{j=255}Im_{i,j}},\rm{and}$ (4a)
$\hat{\phi}_{y}=\frac{\sum_{i=-256}^{i=255}\sum_{j=-256}^{j=255}M_{y_{i,j}}\times Im_{i,j}}{\sum_{i=-256}^{i=255}\sum_{j=-256}^{j=255}Im_{i,j}}.$ (4b)
@@ 5.2 Separation Estimation- Direct Imaging
Because in this experiment we have two point sources separated in 1D, we can make use of our $x$ and $y$ centroid estimates and describe the intensity pattern for direct imaging data on the detector (conditioned on a point source separation of $\theta$) as
${\Psi(x,y|\theta)=\left|\psi\left(x-\hat{\phi}_{x}+\frac{\theta}{2},y-\hat{\phi}_{y}\right)\right|^{2}+\left|\psi\left(x-\hat{\phi}_{x}-\frac{\theta}{2},y-\hat{\phi}_{y}\right)\right|^{2}.}$ (5)
@@
$P_{D}(\{x_{i}\}_{i=1}^{n_{1}}|\theta)=\prod_{i=1}^{n_{1}}\Psi(x_{i},y|\theta),$ (6)
where $x_{i}$ stands for the arrival position of the $i^{\text{th}}$ detected photon, and $n_{1}$ stands for the total number of photons detected during the direct imaging measurement.
For the case when the entire integration time is used for direct imaging, the arrival locations of our photon counting procedure can be used to estimate the parameter of interest. By maximizing the log-likelihood function we get our estimation for the point-source separation:
$\hat{\theta}_{\rm{direct}}=\mathrm{argmax}_{\theta}\left(\log[P_{D}(\{x_{i}\}_{i=1}^{n_{1}}|\theta)]\right).$ (7)
@@ 5.3.1 Maximum Likelihood
In our experiment, we sorted PAD basis modes for the Airy disk PSF. However, for simplicity, and due to the similarity of their plots in the super-resolution regime we matched our binary SPADE photon counting results to the probability of a photon landing in the 0th and 1st Hermite-Gauss modes as a function of the separation, $\theta$, which can be described as such:
$\psi_{0}$ $={e^{-\theta^{2}/\theta_{\mathrm{res}}^{2}}},\quad\mathrm{and}$ (8a)
$\psi_{1}$ $={2\frac{\theta^{2}}{\theta_{\mathrm{res}}^{2}}e^{-\theta^{2}/\theta_{\mathrm{res}}^{2}}}\approx 1-\psi_{0},$ (8b)
The $\psi_{1}$ function can be approximated as it is in Eq. 8b in order to model a binary SPADE, such that all the photons that will not be counted in the $\psi_{0}$ spot, will be counted from the $\psi_{1}$ spot.
@@
$P_{B}(I_{0},n_{2}|\theta)=\begin{pmatrix}n_{2}\\ I_{0}\end{pmatrix}\psi_{0}^{I_{0}}\psi_{1}^{n_{2}-I_{0}},$ (9)
@@
Now, to estimate the most likely $\theta$ we maximize the log-likelihood function $P_{B}$:
$\hat{\theta}_{\rm{Sorting}}=\mathrm{argmax}_{\theta}\left(\log[P_{B}(I_{0},n_{2}|\theta)]\right).$ (10)
@@ 5.3.2 Comparison to Characterization
Another separation estimation method, we used is a direct comparison between the photon numbers detected in each mode and the expected relative intensity shown in 3. The exact number of photons that gets detected at every data point will vary. Because of that, rather than comparing the number of photons in each mode to our characterization we found the ratio between $\psi_{0}$ and $\psi_{1}$ from our data and matched it to the same ratio from our interpolated system’s characterization (Fig. 3). By matching the ratios we could then find what is the separation that is the most fitting for the data. Note that this method will give us a discrete, rather than a continuous set of solutions. On the other hand, when using this method, experimental deviations, using our system, away from the theoretical values don’t hurt the estimation.
@@ 5.4 Separation Estimation- 2-Stage
As mentioned in section 3A, when we try to estimate $\theta$ using our mode sorter, we still use direct imaging for about half of the photons for the centroid estimation. In this method we use the 5000 photons in direct imaging for two purposes, instead of just using these photons for centroid estimation we also use them for the separation estimation. This leads to a new likelihood function that utilizes all the photons collected:
$P_{{\operatorname{2-stage}}}(\{x_{i}\}_{i=1}^{n_{1}},I_{0},n_{2}|\theta)=P_{D}(\{x_{i}\}_{i=1}^{n_{1}},\phi_{x}|\theta)\times P_{B}(I_{0},n_{2}|\theta).$ (11)
Therefore, the estimated $\theta$ using this method can be found by maximizing the log-likelihood function for the composite data:
$\hat{\theta}_{{\operatorname{2-stage}}}=\mathrm{argmax}_{\theta}\left(\log[P_{{\operatorname{2-stage}}}(\{x_{i}\}_{i=1}^{n_{1}},I_{0},n_{2}|\theta)]\right).$ (12)
5.5 Statistical performance evaluation
We computed two statistical quantities to summarize the performance of our various receivers. The first is the mean squared error (MSE) of the estimation procedure with respect to the true value of the parameter $\theta$, defined for a given measurement as
$\textrm{MSE}(\hat{\theta}_{\rm meas})=\left\langle\left(\frac{\hat{\theta}_{\rm meas}-\theta}{\theta_{\rm res}}\right)^{2}\right\rangle.$ (13)
The second measure of performance is the variance of the separation estimate, defined by
$\textrm{Var}(\hat{\theta}_{\rm meas})=\left\langle\left(\frac{\hat{\theta}_{\rm meas}}{\theta_{\rm res}}\right)^{2}\right\rangle-\left\langle\frac{\hat{\theta}_{\rm meas}}{\theta_{\rm res}}\right\rangle^{2}.$ (14)
Another relevant summary statistic of the experimental data processing will be the estimator bias, given by (see Supplementary Materials)
$\textrm{B}(\hat{\theta}_{\rm meas})=\left\langle\frac{\hat{\theta}_{\rm meas}-\theta}{\theta_{\rm res}}\right\rangle.$ (15)
Note that the source separation is written in Rayleigh units for all three of these quantities. From these definitions,
$\textrm{MSE}(\hat{\theta}_{\rm meas})=\textrm{Var}(\hat{\theta}_{\rm meas})+\textrm{B}(\hat{\theta}_{\rm meas})^{2}.$ (16)
@@ 6 Results
Figure 6: Monte-Carlo simulation that shows the mean squared error (MSE) of separation estimation using 10000 photons of direct imaging (blue circles) (refer to section 5.2 for further details), 5000 photons in direct imaging for centroid estimation, and 5000 photons for sorting (orange circles) (refer to section 5.3.1 for further details), and 5000 photons in direct imaging for centroid estimation and for separation estimation, and 5000 photons for sorting (yellow circles) (refer to section 5.3.3 for further details).Not shown on the graph, yet important to note is the quantum Cramer-Rao bound for MSE that is given as follows:$\frac{4\theta_{\rm res}^{2}}{N}=6.4904e-08$, where $\sigma$ is defined in equation 1,and $N$ is the total number of photons. Note that because in our simulation we assumed the absence of bias the plot for the MSE vs. the separation will look identical to the plot for variance vs. separation
@@
Figure 7: a) Mean squared error of the separation estimation , b) Variance of the separation estimation using 10000 photons of direct imaging (blue circles) (refer to section 5.5 for further details), 5000 photons in direct imaging for centroid estimation and 5000 photons for sorting with the maximum likelihood method (orange circles) (refer to section 5.3.1 for further details), 5000 photons in direct imaging for centroid estimation and 5000 photons for sorting with the comparison to characterization method (purple circles) (refer to section 5.3.2 for further details), and 5000 photons in direct imaging for centroid estimation and for separation estimation, 5000 photons for sorting with the maximum likelihood method (yellow circles)(refer to section 5.3.3 for further details).
@@
Before conducting the experiment, we ran a Monte Carlo simulation, which replicates the results of [1] with the hard circular aperture, simulating the estimation procedure for direct imaging, mode sorting, and a two-stage receiver with equal time-share between the two. The results of the simulation are promising as the separation estimation MSE in the deep sub-Rayleigh regime (below 0.5 Rayleigh units of separation) shows an improvement by a factor of up to 12.7. Note that in the Monte Carlo simulation, we assumed a perfect mode-sorter, with no crosstalk, (see supplementary information), and a perfect detector (no noise). This simulation comes to show what MSE an ideal mode sorter can achieve. The simulation results can be seen in Fig. 6. In this figure, when looking at the simulation results when only using the sorting data for separation estimation (orange circles), one can notice a counter-intuitive phenomenon, the MSE increases with the separation. This cam be explained by the fact that at very small separations most of the Fisher information exists in the modes that we sort (PAD00 and PAD01) and as the separation increases there is more Fisher information that lives in higher order modes. Another way to think about it is by looking at fig.3. When the separation is very small, there are very few photons arriving at the PAD01 spot and every change in that is easily detected. However, as the separation increases, small changes at the intensity of the PAD01 spot become harder to detect and therefore making an error becomes more likely.
Figs.7a and 7b present our experimental results for point source separation estimation in terms of the estimator MSE and variance, respectively. In a real-world imaging scenario, MSE is the most relevant quantity, as it directly quantifies the average error from the true value of the parameter. For 0.2 Rayleigh units of separation, our two-stage receiver achieves an MSE 2.76 times lower than experimental direct imaging.
However, we do not see a consistent trend of sub-Rayleigh superiority over direct imaging in terms of MSE in Fig. 7. To understand this, we can recall from Eq. (16) that $\textrm{MSE}(\hat{\theta}_{\rm meas})=\textrm{Var}(\hat{\theta}_{\rm meas})+\textrm{B}(\hat{\theta}_{\rm meas})^{2}$. In Section 6 of the Supplementary Material, we show that our mode sorting measurement exhibited significant estimator bias. The weakness in accuracy of our second-stage measurement is a significant limiting factor in achieving advantageous MSE over direct imaging. This biased estimation is most likely a result of systematic errors in our experiment, e.g., unmitigated crosstalk between spatial modes.
On the other hand, Fig. 7b reveals that the estimator variance for our mode sorting measurement (which includes adaptive alignment from the stage-one centroid estimate) is consistently superior to direct imaging below 0.4 Rayleigh units of point source separation. The same results, neglecting the bias, are shown in Fig. 8. In that figure, one can see a more statistically descriptive plot of the separation estimation variance using direct imaging and mode sorting. In this plot, it becomes apparent that at the deep Rayleigh regime, our mode sorter will give more consistent results with a smaller spread of the estimates.
Figure 8: This plot shows the variance of the separation estimation using sorting (yellow) and direct(blue). In this plot, we subtracted the mean of the of the estimation such that all bias is neglected in this plot. On each box, the central mark indicates the median, and the bottom and top edges of the box indicate the 25th and 75th percentiles, respectively. The whiskers extend to the most extreme data points not considered outliers, and the outliers are plotted individually using the ’+’ marker symbol.
Estimator variance is the statistic that quantifies the precision achievable with a particular measurement scheme, and the most common information theoretic measure for parameter estimation, the Fisher information, is used to bound the estimator variance of any unbiased estimator [4] In other words, while our experimental data may be corrupted by systematic errors that produce estimator bias, the superior estimator variance reveals that our experiment succeeds in demonstrating an adaptive measurement scheme with fundamentally reduced random noise compared with direct imaging for sub-diffraction parameter estimation. Under optimized experimental conditions we expect that our informationally advantageous measurement scheme would reliably produce reduced MSEs for sub-Rayleigh imaging. Some possible sources of systematic error that may have impeded our estimation accuracy in this experiment include: crosstalk, although it was taken into account in the data processing, misalignment of the system, generally and with respect to the centroid, and shot noise in the detector (see Supplementary Material Section 6 for further discussion of the sources of error).
When comparing Fig.7a and Fig.7b it is notable that the characterization method (purple) performs better than sorting (orange) in MSE, but worse in variance. We believe that it performs better in MSE since the in situ characterization inherently corrects for systematic errors leading to estimator bias (see Supplementary Materials Section 6). However, the variance is larger in this method due to the discretization of the possible separations that this estimator can choose, reducing estimation precision.
As a final evaluation of the estimation precision achievable with our adaptive receiver, Fig. 9 directly compares the variance of our experimental estimator against that of our Monte Carlo simulation. Specifically, Fig. 9 plots the ratios between the estimator variances of direct imaging and two-stage mode sorting and compares that ratio for our experiment and simulation. This visualization of our data reveals that, while the idealized simulation understandably produces lower errors than our experiment, the factor by which the mode sorting receiver over- or underperforms direct imaging is largely consistent between simulation and experiment. Fig. 9 shows that our experiment demonstrated roughly all of the relative gain in estimation precision over direct imaging that should be expected based on the simulation results.
Figure 9: Ratio of the variance of separation estimation with direct imaging over the variance of separation estimation with mode sorting as a function of separation in our experiment (blue) and in our Monte-Carlo simulation (orange).
7 Conclusion
Multi-plane light conversion (MPLC) involves passing a multi-spatial mode optical field through multiple stages of programmed spatial-phase-modulation interspersed with fixed-length propagation segments in an isotropic medium. With sufficiently many stages, and sufficiently large transverse cross sections of spatial phase modulation (so as to avoid truncation of the propagating field), MPLC allows for the realization of a general unitary transformation on the span of a set of co-propagating orthogonal spatial modes. We built an MPLC where the spatial phase modulation of the field was realized using distinct segments of a single spatial light modulator (SLM). Using this MPLC, we realized a fully-reconfigurable spatial mode sorter, i.e., a device that spatially separates a number of mutually orthogonal spatial mode functions of any specified modal basis, into spatially-separated beams which can then be detected using individual detectots or assigned detector-pixel segments of a pixelated camera. We showed sorting various orthogonal mode sets, including the binary-phase Hadamard mode sets, the 2D Hermite-Gauss modes, and what we call the PSF-adapted (PAD) basis modes corresponding to the hard-circular-aperture pupil of our imaging system. We showed that our reconfigurable mode sorter can sort two orthogonal modes with less than $1\%$ power cross-talk, and can sort four modes with about $5\%$ power cross-talk. Using this mode sorter, we built a partially-emulated experiment to demonstrate improved spatial resolution—i.e., lower-variance estimation of the angular separation between two equally-bright incoherently-radiating quasi-monochromatic point sources with a sub-Rayleigh separation—over using direct detection using a conventional image-plane focal-plane array. Our receiver follows the recipe from Ref. [1] where the receiver starts out with direct detection as the first stage to estimate the centroid of the two-point-source constellation, and switches to a three-element PAD-basis MPLC mode sorter—at half the total allocated integration time—whose center was electronically-programmed, leveraging the programmability of the MPLC, using the (noisy) estimate of the centroid obtained from the first stage. While our experiment made limited use of the reconfigurability of our MPLC mode sorter, the ability to switch between modal bases is applicable in more challenging multi-source localization tasks with more unknown prior information in subsequent investigations [26]. Building an MPLC-assisted super-resolution imaging system that outperforms direct imaging despite the additional losses internal to the MPLC remains an important future direction. Finally, our experiment used a digital-domain incoherent combination of the detection outcomes—both for direct detection as well as for mode sorting—from a single spatially-scanned laser-light source to mimic the intensity pattern. Extending this work to an experiment with a thermal light and a pinhole mask to mimic a true multi-point source scene, remains an important subject of future research.
Funding
DARPA IAMBIC Program under Contract No. HR00112090128.
Acknowledgments